OpenStudy (anonymous):

From a point A that is 8.90 meters above level ground, the angle of elevation of the top of a building is 32°10' and the angle of depression of the base of the building is 14°10'. Approximate the height of the building. (Round your answer to the nearest tenth.)

OpenStudy (campbell_st):

draw a picture  find the length of the dotted line tan(14 10) = 8.9/d d = 8.9/tan(14 10) the dotted line has length 35.2587 metres now use the angle of elevation and the dotted line length to find the height y tan (32 10) = y/35.2587 y = tan(32 10) x 35.2587 y = 22.175 the total height of the building is 22.175 + 8.9 Approx building height is 31.075 metres